OpenStudy (anonymous):

Find the sixth term of an arithmetic sequence with t1 = 2 and tn = tn-1 + 4.

OpenStudy (zzr0ck3r):

what about your other question?

OpenStudy (anonymous):

I figured it out

OpenStudy (zzr0ck3r):

t1 = 4 t2=t1-4 t2=0 t3=t2-4 t3=0-4 t4=t3-4 t4=-4-4=-8

OpenStudy (zzr0ck3r):

medal and close it then:)

OpenStudy (zzr0ck3r):

the other one, not this.